# Collaborative and parametric insurance on the Ethereum blockchain

### v1

@@ 4 Recursive calculation of the solvency capitals
Figure 1: SCR and MCR over time using the normal approximation.
5 Interaction with the smart contract
During its development phase, a smart contract is not immediately deployed on the Ethereum mainnet. Instead, testnets are used to mimic the behavior of the real blockchain, allowing developers to ensure that the smart contract functions as intended. For our project, we deployed the smart contract on the Ethereum testnet called Sepolia. It features a block explorer, known as Etherscan [^5] See https://sepolia.etherscan.io/, which provides an API to retrieve data. Our contract, named ParametricInsurance, is located at the address:
@@ 5 Interaction with the smart contract
It can be viewed on Etherscan666see https://sepolia.etherscan.io/address/0xa4f3a8567e1d712fc8a53b3a1b6ec76dd42cc10b. The Solidity code is thoroughly discussed in the online supplementary materials accompanying this paper [^7] https://github.com/LaGauffre/smart_parametric_insurance/.
The deployment of the contract was carried out using Remix IDE [^8] See https://remix.ethereum.org/. Remix IDE is an open-source, web-based integrated development environment (IDE) designed specifically for writing, testing, and deploying smart contracts on the Ethereum blockchain. The Solidity code was first compiled before being deployed to the blockchain. Once deployed, the contract is associated with an Ethereum address, allowing interaction through Remix IDE by calling functions and checking the values of the state variables.
@@
ETH Address Role Intial balance (in test net ETH)
0xE8e79B8B8c0481fa33a8E0fcA902ad5754BfE1C3 owner $0.4366$
0x2CF8ed1664616483c12Ef3113f6F62E68f1a810A surplus provider $0.2$
0xd34a37613A382bA503f1599F514C9788dF3659C4 policyholder $0.21$
0xa4f3a8567e1d712fc8a53b3a1b6ec76dd42cc10b smart contract 00
@@
We designed a scenario that calls all the functions available in the smart contract described in Section 3. Here, we take an example involving some weather risks (hail, wind, …) for illustration purposes, but other risks could be considered (flight delays, event cancellation, …). The steps are as follows:
@@
    Owner deploys the ParametricInsurance contract with the initial parameters:
    $\eta^{(1)}=0.1,\quad\eta^{(2)}=0.05,\quad\alpha_{\text{SCR}}=0.995,\quad\text{and }\alpha_{\text{MCR}}=0.85.$
@@
    Surplus provider sends ETH 0.1 to the smart contract.
@@
    Owner sends ETH 0.08 to the smart contract.
@@
    Policyholder underwrites policy #1 with characteristics ("$rain",0.06\right)$, where:
    $\theta_{1}=\mathbb{P}("rain")=0.7164.$
    Recall from Remark 4.1 that underwriting is done by providing a description of the event as a character string. The function getRandomNumber returns a probability for the event, which is compared to the probability obtained by applying the same function to the description of the observed event.
@@
    Policyholder underwrites policy #2 with characteristics ("$hail",0.01\right)$, where:
    $\theta_{2}=\mathbb{P}("hail")=0.0315.$
  6. 6.
    Owner updates the parameters:
    $\eta^{(1)}=0.025,\quad\eta^{(2)}=0.025.$
  7. 7.
    Policyholder underwrites policy #3 with characteristics ("$snow",0.1\right)$, where:
    $\theta_{3}=\mathbb{P}("snow")=0.5328.$
  8. 8.
    Owner settles policy #2 with "hail1" as the observed event, where:
    ("$\mathbb{P}("hail1")=0.4496>\theta_{2},$
    resulting in no compensation.
  9. 9.
    Owner burns 0.1 protocol tokens.
  10. 10.
    Owner settles policy #3 with "snow2" as the observed event, where:
    ("$\mathbb{P}("snow2")=0.2058<\theta_{3},$
    triggering a compensation of $l_{3}=\text{ETH}0.1$.
  11. 11.
    Policyholder underwrites policy #4 with characteristics ("$sun",0.08\right)$, where:
    $\theta_{4}=\mathbb{P}("sun")=0.4188.$
  12. 12.
    The Surplus provider burns $0.0157$ protocol tokens, causing the surplus to reach the SCR level. In other words, the Surplus provider withdraws the maximum amount allowable.
  13. 13.
    Owner changes the parameter:
    $\alpha_{\text{MCR}}=0.985.$
  14. 14.
    Owner settles policy #4 with "sun1" as the observed event, where:
    ("$\mathbb{P}("sun1")=0.0128<\theta_{4},$
    triggering a compensation of $l_{4}=\text{ETH}0.08$.
  15. 15.
    Policyholder underwrites policy #5 with characteristics ("$wind",0.02\right)$, where:
    $\theta_{5}=\mathbb{P}("wind")=0.4835.$
  16. 16.
    Owner changes the parameter:
    $\alpha_{\text{MCR}}=0.99.$
  17. 17.
    Owner burns $0.023$ protocol tokens, causing the surplus to reach the SCR level. In other words, the Owner withdraws the maximum amount allowable.
  18. 18.
    Owner settles policy #5 with "wind9" as the observed event, where:
    ("$\mathbb{P}("wind9")=0.1940<\theta_{5},$
    triggering a compensation of $l_{5}=\text{ETH}0.02$.
The final event causes the smart contract to file for bankruptcy, as the surplus $X$ falls below the MCR threshold. Consequently, the premium for policy #$\#1$ is refunded to the policyholder. The remaining surplus is then distributed among all token holders, including the owner, surplus provider, and the policyholder.
We do not claim that the participants in this scenario acted strategically—in fact, quite the opposite. By withdrawing the maximum amount allowable on two occasions, both the Surplus provider and the Owner significantly jeopardized the solvency of the smart contract. Additionally, the parameter adjustments, which made the SCR and MCR nearly identical, are not representative of realistic scenarios. Lastly, the situation where multiple insurance contracts result in compensation is both highly unlucky and statistically improbable. These choices were intentional, designed to illustrate the consequences of the smart contract going bankrupt and undergoing a reset.
Deploying the contract results in a "contract creation" transaction on the blockchain. Each subsequent function call also generates a transaction recorded on the blockchain. These transactions can be retrieved and organized into a data frame, the structure of which is detailed in Table 2.
@@
blockNumber int 7143345 Number of the block in the blockchain
blockHash hex 0xf8d9aa8$\ldots$ Hash value of the block
timeStamp int 1732457004 Time at which the transaction got recorded
hash hex 0x93cf64$\ldots$ Hash value of the transaction, servees as identifier on the blockchain
transactionIndex int 10 Location of the transaction in the block
from hex 0xe8e7$\ldots$ Address of the transaction sender
to hex Address of the transaction recipinet, here it is the smart contract address
value int 0 The amount transferred within the transaction
gas int 2935995 gas limit to be spend by the transaction sender
gasPrice int 12145467087 Value of the unit of gas (in WEI where $\text{WEI}1=\text{ETH}10^{-18})$
input data 0x6080604052612$\ldots$ The input data of tthe transaction
methodId hex 0x60806040$\ldots$ Identifier of the function being called
functionName string "Fund" Name of the function being called
contractAddress hex 0xa4f3a856$\ldots$ Address of the smart contract
gasUsed int 2911719 The actual amount of gas used to send the transaction.
@@
The value of a unit of gas (gasPrice) fluctuates over time, as it depends on the level of activity on the Ethereum network. High network congestion leads to higher gas prices. Another factor influencing transaction costs is the amount of data being recorded, which is reflected in the input field of the transaction details.
@@
ParametersUpdated Update newEta1 int $1000$ premium loading
newEta2 int $500$
@@
customer hex $0x456\ldots$ Ethereum address
EventDescription string "rain" Description of the insured event
payoutAmount int 1000000 Compensation if the event occurs
@@
customer hex $0x456\ldots$ Ethereum address
@@
Xt int 8000000 Surplus
@@
The events used in this study are defined in the Solidity code, which is included in the online supplementary materials accompanying this article. Using Python, we retrieve the event logs to track participant balances over time. Figure 3 shows the balances (in ETH) of the Ethereum addresses associated with the Owner, the Surplus Provider, the Policyholder, and the Smart Contract after each call to the function.
@@
We observe that the scenario was particularly beneficial to the Policyholder, who received three compensation payments and a refund for the first policy she purchased. The event logs also enable us to visualize the trajectories of the various processes defined in Section 3. Specifically, Figure 4 displays the trajectories of $X$ (surplus), $B$ (balance), SCR (Solvency Capital Requirement), and MCR (Minimum Capital Requirement) as they evolved throughout the scenario.
@@ 6 Limitations and perspectives
This article proposes a framework for a simplified smart contract in insurance and demonstrates the feasibility of its implementation on a blockchain. Additionally, it makes this open-source prototype available to other researchers, facilitating further exploration of its limitations and potential enhancements. In this section, we outline the main limitations of this initial approach and suggest possible solutions.
The decentralized nature of the contract and the parametric nature of its trigger offer significant cost-saving opportunities in claim management, distribution, and fraud prevention. For example, in general insurance contracts, fraud—though inherently difficult to quantify—may account for as much as $5\%$ to $15\%$ of collected premiums. In France, insurance fraud was estimated at a minimum of 466 million euros in 2023, representing a $50\%$ increase from 2022101010Source: ALFA (https://www.alfa.asso.fr/), the french Insurance Fraud Control Agency.
Some features that are simplified in this prototype (as complexity increases costs) could be improved without compromising the smart contract’s attractiveness in terms of pricing. However, addressing these improvements must account for potential challenges, such as reputation risks or other factors that might discourage customer participation. Enhancing risk computations, compliance, and risk management must be carefully balanced. Excessive complexity risks becoming counterproductive, akin to the trade-offs observed in the ruin prevention framework discussed by Gauchon et al. [14] and Gauchon et al. [15].
@@
In our simplified approach, claims are assumed to be independent, and the underlying risks are considered stationary. However, in practice, many risks suitable for smart contracts with parametric triggers are influenced by regime changes, which can lead to event clustering or increased claim frequency. To mitigate adverse selection, a critical first step in risk management is introducing a waiting period before the guarantee becomes active. This measure helps prevent policyholders from exploiting arbitrage opportunities when they are aware of worsening risks. For instance, in the case of flight disruptions, events like strikes, conflicts, or volcanic eruptions can make delays almost certain in the short term. Similarly, for weather-related risks, extreme event warnings may be issued days in advance. It is essential to ensure that policyholders cannot exploit the smart contract under such circumstances while paying the same premium as usual.
In some cases, a waiting period may not be sufficient to mitigate risk, as underlying risks can be anticipated several months in advance. For example, the El Ni no/La Ni na cycle, which is partially predictable months ahead, increases the likelihood of certain extreme weather events while reducing others. Moreover, long-term factors such as climate change and sectoral inflation can exacerbate risks over time. If competitors actively adjust their premiums to reflect these changes while the smart contract does not, adverse selection becomes a likely outcome.
To address these challenges, the model should incorporate additional complexity, such as accounting for correlations. For instance, Denuit and Robert [9] explore the use of Ising models in collaborative insurance to capture such dependencies. Additionally, finding a way to update premiums dynamically is essential. This could be achieved either automatically—by scraping new data from competitors, insurance federations, or other relevant sources—or through decisions made by a "board" of customers and investors, who could leverage voting rights granted via tokens. While these refinements would increase costs, they seem practical given the potential savings on fraud and claim management.
The absence of staff turnover risk and the near-instantaneous resolution of claims offer significant advantages over traditional insurance models. High turnover in claims management teams, for instance, can lead to critical lapses in detecting changes in claim frequency.
In addition to systemic (market-wide) changes, insurers may encounter higher-than-expected claim frequencies. Adapting premiums using experience rating and credibility theory could help counteract potential adverse selection, provided the associated costs remain manageable. To complement this approach, the risk capital module should also be updated to reflect these credibility adjustments. This could involve leveraging large deviation principles rather than relying solely on central limit theorems, as discussed in Trufin and Loisel [25].
@@
In both cases, it is worthwhile to establish a governance system where token-holding investors form a board with voting rights to decide on key parameters such as the safety loading level, premium principles, and the risk level $\alpha$. This aligns with the concept of decentralized organizations, which are extensively studied in management science and have naturally emerged within blockchain applications. However, classical issues such as over- or under-reactions stemming from individual risk perceptions and cognitive biases could be amplified, especially if voter participation is low or voting power is concentrated due to coalitions. These challenges are not uncommon, as low or uneven participation rates have frequently been observed in various decentralized networks, we refer the reader to the work of Messias et al. [20] for a comprehensive discussion.
As policyholders receive tokens when they underwrite a contract, they effectively become investors, creating a collaborative insurance framework. In this context, premium calculation could be based on the conditional mean risk-sharing principle, as proposed by Denuit and Robert [10].
@@
Our smart contract is exposed to the risk of a "mass investor lapse," where investors may collectively withdraw their tokens. Stronger incentive mechanisms could be implemented to encourage investors to retain their tokens, particularly when the contract’s balance approaches the risk capital threshold. One potential approach is to introduce a coin age principle, where tokens held for longer periods gain additional value. This concept has already been explored in the implementation of the cryptocurrency PeerCoin [^11] See https://www.peercoin.net/. A traditional approach to managing the surplus of an insurance company is to implement a dividend policy that rewards investors or, more generally, token holders. This can be achieved using classical optimal dividend strategies. Barrier or band strategies, for instance, have been shown to be optimal across various modeling frameworks (see the survey by Avanzi [3]). In our framework, we operate within a discrete-time risk model with capital injections and withdrawals. This approach aligns with the structure of our system, where wealth is only inspected at discrete points—namely, during settlements or when an investor funds or withdraws from the contract.
@@
In our approach, surplus management is handled through proxies referred to as the Minimum Capital Requirement (MCR) and Solvency Capital Requirement (SCR). If such a smart contract were to be classified as an insurance product, it would need to comply with the absolute minimum MCR threshold prescribed by Solvency II in Europe. This requirement could lead to a chicken-and-egg problem, as achieving the necessary critical mass of funds would be essential to meet the MCR. Additionally, calculating the SCR and MCR would typically involve integrating multiple risk modules, adding further complexity. Enhancing the simplified computations of these capital requirements would also incur additional implementation costs.
If the contract is not classified as an insurance product, it could lead to tax and regulatory implications. To ensure recognition as insurance, it is crucial to keep basis risk sufficiently low. Effective basis risk management not only protects policyholders but also mitigates the risk of the contract being reclassified as gambling. In some countries, policyholders may be required to provide brief proof of loss to validate claims. While we assume that the occurrence of a compensation-triggering event inherently results in a loss for the policyholder, real-world scenarios might necessitate additional efforts from either the policyholder or the smart contract’s governance system to address such requirements.

### v2

@@ 5 Interaction with the smart contract
It can be viewed on Etherscan555see https://sepolia.etherscan.io/address/0xbe035cf1367c45A0C9517969F5ABDd3abF743ae7. The complexity of the insurance mechanism necessitated the deployment of two additional smart contracts. PricingLogic located at the address
@@
calculates the premium and ModelPointsLogic located at
[…]
computes the solvency capitals. The Solidity code for these smart contracts is thoroughly discussed in the online supplementary materials that accompany this paper. [^6] https://github.com/LaGauffre/smart_parametric_insurance/.
The deployment of the contract was carried out using Remix IDE [^7] See https://remix.ethereum.org/ which is an open-source, web-based integrated development environment specifically designed for writing, testing, and deploying smart contracts on the Ethereum blockchain. We first compile the Solidity code before deploying it to the blockchain. Once deployed, the contract is associated with an Ethereum address, enabling interaction through Remix IDE by calling functions and checking the values of the state variables.
@@
ETH Address Role
0xE8e79B8B8c0481fa33a8E0fcA902ad5754BfE1C3 owner
0x2CF8ed1664616483c12Ef3113f6F62E68f1a810A surplus provider
0xd34a37613A382bA503f1599F514C9788dF3659C4 policyholder
0xbe035cf1367c45A0C9517969F5ABDd3abF743ae7 smart contract
@@
Our smart contract allows for the underwriting of parametric insurance contracts to compensate policyholders if the height of precipitation exceeds a threshold $\bar{Q}=5$ on a particular day in the year 2025 in either Marseille or Strasbourg. The Cornish-Fisher approximation is employed as soon as we reach 5 model points. The day of the year is represented as an integer between 1 and 365. We designed the following scenario to illustrate the functions available in the smart contract described in Section 4.
@@
    Deployment: The Owner initiates the process by deploying three key contracts: PricingLogic, ModelPointsLogic, and InsuranceLogic. The initial parameters are set as follows:
    $\eta=0.1,\quad\alpha_{\text{SCR}}=0.995,\quad\text{and}\quad\alpha_{\text{MCR}}=0.85.$
@@
    Initial Funding: A Surplus Provider kicks things off by sending 0.1 ETH to the smart contract, providing the initial capital needed to underwrite policies.
@@
    Policy Underwriting: Policyholders begin underwriting various insurance policies to protect against precipitation exceeding 5 units in Marseille or Strasbourg on specific days of the year. Here are the details of the policies they secure:
    • Policy #1: Covers Marseille on day 60, with a liability of $0.01$ ETH.
    • Policy #2: Covers Strasbourg on day 60, with a liability of $0.01$ ETH.
    • Policy #3: Covers Marseille on day 45, with a higher liability of $0.02$ ETH.
    • Policy #4: Covers Strasbourg on day 80, with a liability of $0.005$ ETH.
    • Policy #5: Covers Strasbourg on day 102, with a liability of $0.015$ ETH.
    • Policy #6: Covers Strasbourg on day 206, with a liability of $0.01$ ETH.
    • Policy #7: Similar to Policy #6, covering Strasbourg on day 206, with a liability of $0.01$ ETH.
    • Policy #8: Covers Marseille on day 300, with a liability of $0.015$ ETH.
    • Policy #9: Covers Marseille on day 282, with a liability of $0.005$ ETH.
    • Policy #10: Covers Marseille on day 180, with a liability of $0.02$ ETH.
@@
    Settling one Contract and withdrawing of the investor: The Owner starts settling the contracts based on the observed precipitation:
    • Contract #3 is settled with precipitation in Marseille on day $45$ not exceeding the threshold.
    • A Surplus Provider decides to burn $0.035$ tokens, adjusting their investment.
@@
    Further Contract Settlements: The Owner continues settling more contracts:
    • Contract #1 is settled with precipitation in Marseille on day $60$ exceeding the threshold.
    • Contract #2 is settled with precipitation in Strasbourg on day $60$ exceeding the threshold.
    • Contract #4 is settled with precipitation in Strasbourg on day $80$ exceeding the threshold.
    • Contract #5 is settled with precipitation in Strasbourg on day $102$ not exceeding the threshold.
The final event triggers the smart contract to file for bankruptcy, as the surplus $X$ falls below the Minimum Capital Requirement (MCR) threshold. Consequently, the premiums for policies #$\#6$, #$\#7$, #$\#8$, #$\#9$, and #$\#10$ are refunded to the policyholders. The remaining surplus is then returned to the surplus provider.
We do not suggest that the participants in this scenario acted strategically—in fact, quite the opposite. By withdrawing funds early, the surplus provider jeopardized the solvency of the smart contract. The scenario where multiple insurance contracts result in compensation is both highly unlucky and statistically improbable. Additionally, the reduction in the number of model points below the minimum required for the Cornish-Fisher approximation led to a drastic increase in the MCR, which ultimately exceeded the surplus. These choices were intentional, designed to illustrate the consequences of the smart contract going bankrupt and undergoing a reset.
Deploying the contract results in a “contract creation” transaction on the blockchain. Each subsequent function call also generates a transaction, which is recorded on the blockchain. These transactions can be retrieved and organized into a data frame, the structure of which is detailed in Table 4.
@@
blockNumber object 8518222 The height of the block in the blockchain
timeStamp object 1749558720 The date and time when the transaction was created or mined
hash object 0xe4$\ldots$ The unique hash value identifying the transaction
nonce object 101 A number used to ensure that each transaction can only be processed once
blockHash object 0x0248$\ldots$ The unique hash value identifying the block that contains the transaction
transactionIndex object 31 The position of the transaction within the block
from object 0xe8e7$\ldots$ The address of the sender who initiated the transaction
to object 0xbe03$\ldots$ The address of the recipient of the transaction
value object 0 The amount of ETH sent with the transaction
gas object 3958691 The maximum amount of gas allocated for the transaction
gasPrice object 1501121993 The price of each unit of gas, specified in ETH
isError object 0 Indicates whether an error occurred during the transaction
txreceipt_status object 1 Indicates the status of the transaction receipt
input object 0x6080$\ldots$ The data field containing information necessary for the transaction
contractAddress object 0xbe03$\ldots$ The address of the contract created as a result of the transaction
cumulativeGasUsed object 6358722 The total amount of gas used in the block up to and including this transaction
gasUsed object 3926362 The amount of gas actually used by this specific transaction
confirmations object 652 The number of blocks mined after the block containing this transaction
methodId object 0x6080$\ldots$ The identifier for the method being called in a smart contract
functionName object The name of the function called in a smart contract
@@
ParametersUpdated Update newEta int $1000$ premium loading
@@
customer__\__address hex $0x456\ldots$ Ethereum address
T int 80 Day of the event with year 2025
station string STRASBOURG-ENTZHEIM location of the event
l int 1000000 Compensation if the event occurs
cp int 1000000 Commercial premium
@@
SCR int 100000 Updated Solvency Capital Requirement
MCR int 100000 Updated Minimum Capital Requirement
@@
customer__\__address hex $0x456\ldots$ Ethereum address
@@
SCR int 100000 Updated Solvency Capital Requirement
MCR int 100000 Updated Minimum Capital Requirement
@@
The events used in this study are defined in the Solidity code, which is included in the online supplementary materials accompanying this article. Using Python, we retrieve the event logs to track participant balances over time. Figure 9 illustrates the balances, in ETH and before transaction fees, of the Ethereum addresses associated with the Owner, the Surplus Provider, the Policyholder, and the Smart Contract after each function call.
@@
We observe that the scenario was particularly beneficial to the Policyholder, who received three compensation payments and a refund for five policies. The event logs also enable us to visualize the trajectories of the various processes defined in Section 3 and Section 4. Specifically, Figure 10 displays the trajectories of $X$ (surplus), $B$ (balance), SCR (Solvency Capital Requirement), and MCR (Minimum Capital Requirement) as they evolved throughout the scenario.
@@
Note the difference between the surplus and the balance of the smart contract: the balance of the smart contract increases at each "Underwrite" event, while the surplus remains constant. Until contract #$\#5$ is underwritten, the Solvency Capital Requirement (SCR) and Minimum Capital Requirement (MCR) are exactly equal to the sum of all potential compensations to be paid. Once contract #$\#5$ is underwritten, the number of model points reaches $5$, and the solvency capitals are calculated via the Cornish-Fisher approximation, causing a drop in the value of the solvency capitals. After the settlement of contract #$\#3$, the exchange rate of protocol tokens against ETH appreciates, which is why the surplus provider decides to cash out part of her investment. The compensations paid after the settlements of contracts #$\#1$, #$\#2$, and #$\#4$, combined with the fact that the settlement of contract #$\#5$ reduces the number of model points below $5$, cause the smart contract to reset as the surplus becomes smaller than the MCR. The status of all active policies changes to "cancelled." Premiums are refunded to the policyholders, and investors receive any remaining balance from the smart contract.
Remark 5.1.
The drop in value of both the Minimum Capital Requirement (MCR) and the Solvency Capital Requirement (SCR) after including contract ##\##5 might seem surprising at first glance. Under Solvency II regulations, there is an absolute floor on the MCR value, set at least to €2.7 million euros for non-life insurance companies. We have proposed an alternative method for defining this floor, based on the sum of exposures when the number of independent risks is too small (fewer than 5 in Figure 10 for illustration purposes, but typically fewer than hundreds or thousands in real-life scenarios). In practice, a decrease in solvency capital can occur during the transition from the standard formula to internal models, which somewhat corresponds to adopting normal approximations in our example.
@@ 6 Limitations and perspectives
This article proposes a framework for a smart contract suited to parametric insurance and demonstrates the feasibility of its implementation on a blockchain. Additionally, it makes this open-source prototype available to other researchers, facilitating further exploration of its limitations and potential enhancements. In this section, we outline the main limitations of this initial approach and suggest possible solutions.
@@
The underlying risks are considered stationary. However, in practice, many risks suitable for smart contracts with parametric triggers are influenced by regime changes, which can lead to event clustering or increased claim frequency. To mitigate adverse selection, a critical first step in risk management is introducing a waiting period before the guarantee becomes active. This measure helps prevent policyholders from exploiting arbitrage opportunities when they are aware of worsening risks. For instance, in the case of flight disruptions, events like strikes, conflicts, or volcanic eruptions can make delays almost certain in the short term. Similarly, for weather-related risks, extreme event warnings may be issued days in advance. It is essential to ensure that policyholders cannot exploit the smart contract under such circumstances while paying the same premium as usual.
@@
In some cases, a waiting period may not be sufficient to mitigate risk, as underlying risks can be anticipated several months in advance. For example, the El Niño/La Niña cycle, which is partially predictable months ahead, increases the likelihood of certain extreme weather events while reducing others. Moreover, long-term factors such as climate change and sectoral inflation can exacerbate risks over time. If competitors actively adjust their premiums to reflect these changes while the smart contract does not, adverse selection becomes a likely outcome. Finding a way to update premiums dynamically is essential. This could be achieved either automatically—by scraping new data from competitors, insurance federations, or other relevant sources—or through decisions made by a "board" of customers and investors, who could leverage voting rights granted via tokens. A governance system where token-holding investors form a board with voting rights aligns with the concept of decentralized organizations, which are extensively studied in management science and have naturally emerged within blockchain applications. However, classical issues such as over- or under-reactions stemming from individual risk perceptions and cognitive biases could be amplified, especially if voter participation is low or voting power is concentrated due to coalitions. These challenges are not uncommon, as low or uneven participation rates have frequently been observed in various decentralized networks, we refer the reader to the work of Messias et al. [22] for a comprehensive discussion.
@@
Our smart contract is exposed to the risk of a "mass investor lapse," where investors may collectively withdraw their tokens. Stronger incentive mechanisms could be implemented to encourage investors to retain their tokens, particularly when the contract’s balance approaches the risk capital threshold. One potential approach is to introduce a coin age principle, where tokens held for longer periods gain additional value. This concept has already been explored in the implementation of the cryptocurrency PeerCoin [^9] See https://www.peercoin.net/. A traditional approach to managing the surplus of an insurance company is to implement a dividend policy that rewards investors or, more generally, token holders. This can be achieved using classical optimal dividend strategies. Barrier or band strategies, for instance, have been shown to be optimal across various modeling frameworks (see the survey by Avanzi [3]). In our framework, we operate within a discrete-time risk model with capital injections and withdrawals. This approach is consistent with the structure of our system, where wealth is only inspected at discrete points—namely, during settlements or when an investor funds or withdraws from the contract.
@@
In our approach, surplus management is handled through proxies referred to as the Minimum Capital Requirement (MCR) and Solvency Capital Requirement (SCR). If such a smart contract were to be classified as an insurance product, it would need to comply with the absolute minimum MCR threshold prescribed by Solvency II in Europe. This requirement could lead to a chicken-and-egg problem, as achieving the necessary critical mass of funds would be essential to meet the MCR. Additionally, calculating the SCR and MCR would typically involve integrating multiple risk modules, adding further complexity and thus incur additional implementation costs.
Exchange rate risk is a significant concern for potential customers, particularly as compensations are paid in ETH, a cryptocurrency subject to high volatility against fiat currencies. To mitigate this issue, one potential solution is to link our contract with another smart contract designed to exchange ETH for a stablecoin. This approach leverages the interoperability of blockchain applications, a key advantage of decentralized finance over traditional financial systems.
If the contract is not classified as an insurance product, it could lead to tax and regulatory implications. To ensure recognition as insurance, it is crucial to keep basis risk sufficiently low. Effective basis risk management not only protects policyholders but also mitigates the risk of the contract being reclassified as gambling. In some countries, policyholders may be required to provide brief proof of loss to validate claims. While we assume that the occurrence of a compensation-triggering event inherently results in a loss for the policyholder, real-world scenarios might necessitate additional efforts from either the policyholder or the smart contract’s governance system to address such requirements. Basis risk management is a central problem when dealing with parametric insurance solution and as such under investigation in many recent research work like that of Niakh et al. [26].